# Rail-only: A Low-Cost High-Performance Network for Training LLMs with Trillion Parameters

### v4

@@ Abstract
This paper presents a low-cost network architecture for training large language models (LLMs) at hyperscale. We study the optimal parallelization strategy of LLMs and propose a novel datacenter network design tailored to LLM’s unique communication pattern. We show that LLM training generates sparse communication patterns in the network and, therefore, does not require any-to-any full-bisection network to complete efficiently. As a result, our design eliminates the spine layer in traditional GPU clusters. We name this design a Rail-only network and demonstrate that it achieves the same training performance while reducing the network cost by 38% to 77% and network power consumption by 37% to 75% compared to a conventional GPU datacenter. Our architecture also supports Mixture-of-Expert (MoE) models with all-to-all communication through forwarding, with only 4.1% to 5.6% completion time overhead for all-to-all traffic. We study the failure robustness of Rail-only networks and provide insights into the performance impact of different network and training parameters.
@@ I Introduction
We evaluate the performance of our Rail-only network architecture using an analytical formulation and provide insights into the performance impact of different network and training parameters. We compare the cost and power consumption of a Rail-only network to a full-bisection bandwidth any-to-any Clos network and show that our LLM-centric network architecture reduces the network cost and power by 38%–76% and 37%–75%, respectively (§V-E). Moreover, we show that a Rail-only network achieves the same performance as a Rail-optimized cluster for LLMs without MoE layers. Finally, we demonstrate that a Rail-only interconnect only incurs 4.1%–5.6% throughput overhead for LLMs with MoEs that require all-to-all traffic.
@@ II-A Intra-platform Connectivity: High-bandwidth Domain
For instance, Nvidia’s DGX H100 server [14] consists of eight H100 GPUs interconnected with NVSwitches, providing 7.2 Tbps of non-blocking bandwidth internally. The GB200 NVL72 computer announced recently connects 36 GB200 Superchips with fifth-generation NVLink within a rack at 14.4 Tbps per GPU [15]. The AMD MI300X platform, on the other hand, employs AMD’s Infinity Fabric to connect eight MI300X accelerators in a full-mesh topology with 7.2 Tbps of bandwidth per GPU [6]. Similar platforms such as Nvidia’s DGX GH200 Super Computer have utilized multi-tiered NVSwich topologies to scale the platform’s size up to 256 GPUs while maintaining 7.2 Tbps full-bisection intra-GPU bandwidth [16]. This paper refers to a platform with Tbps internal bandwidth connectivity as a “high-bandwidth (HB) domain”, and the corresponding interconnect as HB interconnects (HBI).
@@ III-B Traffic in the NIC Domain for LLMs
Figure 2a shows the traffic pattern of training GPT-1T with hierarchical collective communication spanning both NIC and HB domains, differing from the MegatronLM cases. We compute and analyze the traffic pattern of training the GPT-1T model, with a batch size of 4096, distributed in a cluster composed of 16 Nvidia DGX GH200 supercomputers [16] (4096 GPUs). Each DGX GH200 supercomputer comprises a two-tier NVSwitch architecture, facilitating 7.2 Tbps GPU-to-GPU bandwidth across 256 H100 GPUs. Additionally, each GPU has a Connect-X7 HCA Infiniband network interface [16], which provides 400 Gbps network bandwidth in/out of each GPU. In this setup, each DGX GH200 supercomputer forms an HB domain. Figure 4 illustrates the traffic volume percentage and heatmap in this setting. The parallelization strategy has a total data parallel degree of 64, which spans 32 GPUs in each HB domain and two HB domains across the network. Figures 2b and 2c illustrate the traffic heatmap of the hierarchical AllReduce algorithm, which splits the AllReduce traffic among each DP group. Note that the network traffic stays within a rail (GPUs with a distance of 256 apart). The hierarchical algorithm efficiently utilized the bandwidth in the HB domain to carry $98\%$ of the AllReduce traffic, while the network domain carries the other $2\%$.
@@ IV-A Architecture Design
Figure 6 illustrates our Rail-only network architecture. Compared to a conventional Rail-optimized GPU cluster, shown in Figure 1, our network keeps the HB domains and rail switches but omits the spine switches. Hence, a Rail-only network provides network connectivity only within the same rail, without changing the bandwidth across GPU pairs.
Figure 6: Our proposal: remove the spine layer to form a Rail-only connection.
A straightforward way to realize our proposed architecture is to remove the spine switches from Figure 1 and re-purpose all the uplinks connecting rail switches to the spine as down-links to GPUs. Hence, a dedicated but separate Clos network connects each rail. Since more ports of the switches in the rail are available as down-links to GPUs, compared to the Rail-optimized network, the Rail-only design saves the number of switches in the network by both removing the spine switches and reducing the number of switches needed per rail. In the rest of this paper, we base our analysis on this realization of the Rail-only network.
@@ IV-B Routing in Rail-only Networks
The two terms differ by ${y(x-1)D}/{C_{F}}$, which is the cost of forwarding within HB domains. When $y(x-1)\approx x(y-1)$, the slow-down factor is approximately $C_{S}/C_{F}$, which equals to $4.1\%$ and $5.6\%$ for the DGX A100 and DGX H100 generations of GPU platforms, respectively. This factor is low because of the bandwidth asymmetry between the two domains. Furthermore, this slow-down only applies to the all-to-all communication, which accounts for $27\%$ of the total traffic as shown in Figure 5. Therefore, this small overhead is negligible by Amdah’s law. We note that such properties also make our network design suitable for other classes of DNN models.
@@ V-A Iteration Time Modeling
Figure 7 compares the HFUs approximated by our analytical model with the ground truth for different GPT models and cluster scales. For GPT-1T, our computed HFU only differs from the ground truth by $0.15\%$. The discrepancy between our analytical model and the ground truth comes from our idealistic modeling of GPU computation and communication, assumptions on how computation and communication overlap, and underestimation of memory overhead. Such discrepancy goes up as the model size decreases. For the GPT-22B model, our estimation error is $15.7\%$ compared to ground truth MFU. The rest of this section utilizes our analytical model to estimate the training iteration time of Rail-only interconnects.
@@ V-B What is the Optimal Size of an HB domain?
As depicted in Figure 8, iteration times decrease as the HB domain size increases, indicating that larger HB domains reduce the network overhead during training. In both GPT models, the iteration time achieved with an HB domain size of 256 is nearly optimal. Compared to the ideal case (all GPUs are under a monolithic HB domain), GPT-146B with an HB domain of 256 is $8.9\%$ slower, while GPT-1T is $1.3\%$ slower. However, the marginal gain decreases as the HB domain size increases. For the larger GPT-1T model, the train iteration time plateaus above 32 GPUs due to Amdhahl’s law, suggesting diminishing returns from the increased cost of bigger HB domains.
@@
For a smaller GPT-146B model, shown in Figure 9b, the marginal gain of increasing HB domain size is higher than that of GPT-1T. Providing an HB domain of size eight reduces the iteration time by $50.6\%$ compared to the HB domain of size one, while increasing the HB domain size from 8 to 256 further achieves a $39.1\%$ reduction in iteration time. The more significant marginal gain for smaller LLMs incurs more communication overhead when distributed to the same cluster than larger models. This effect arises from how computation and communication costs scale as LLM grows. The communication requirement increases linearly with the model’s hidden size and sequence length. On the other hand, the model FLOPs increase quadratically with these two parameters, as indicated by previous work [21]. Therefore, we conclude that the optimal HB domain size depends on the size of the GPT model.
@@ V-D Impact of Batch Size on Network Design
While the batch size is typically an ML-centric metric for optimization, our analysis indicates that the impact of batch size on the total training time goes beyond the total number of iterations required for convergence. To further understand such impact, we analyze the iteration time of a GPT-1T model on a 32768 GPU cluster while changing the HB domain size from $K=8$ to $32768$. In this study, we vary the global batch size from 256 to 4096. Figure 10a plots the change in iteration time as the batch size varies. The iteration time exhibits a similar trajectory for all HB domain sizes; however, the relative performance (the ratio of the iteration time for an HB domain size to that of the ideal case) improves as the batch size increases. Figure 10b represents this trend. When $K=256$, the relative performance increases from $93\%$ to $99\%$ as the batch size increases from 256 to 4096 sequences. The iteration time advantage is prominent when the HB domain size is small. For $K=8$, increasing the batch size from 256 to 4096 improves the relative performance from $58\%$ to $82\%$, suggesting a larger batch size is preferable for a cluster with a smaller HB domain. Prior studies have shown that LLM training benefits from a larger batch size [29, 1], especially for bigger models, making it a perfect fit for our Rail-only design.

### v5

@@ Abstract
This paper presents a low-cost network architecture for training large language models (LLMs) at hyperscale. We study the optimal parallelization strategy of LLMs and propose a novel datacenter network design tailored to LLM’s unique communication pattern. We show that LLM training generates sparse communication patterns in the network and, therefore, does not require any-to-any full-bisection network to complete efficiently. As a result, our design eliminates the spine layer in traditional GPU clusters. We name this design a Rail-only network and demonstrate that it achieves the same training performance while reducing the network cost by 38% to 77% and network power consumption by 37% to 75% compared to a conventional GPU datacenter. Our architecture also supports Mixture-of-Expert (MoE) models with all-to-all communication through forwarding, with only 8.2% to 11.2% completion time overhead for all-to-all traffic. We study the failure robustness of Rail-only networks and provide insights into the performance impact of different network and training parameters.
@@ I Introduction
We evaluate the performance of our Rail-only network architecture using an analytical formulation and provide insights into the performance impact of different network and training parameters. We compare the cost and power consumption of a Rail-only network to a full-bisection bandwidth any-to-any Clos network and show that our LLM-centric network architecture reduces the network cost and power by 38%–76% and 37%–75%, respectively (§V-E). Moreover, we show that a Rail-only network achieves the same performance as a Rail-optimized cluster for LLMs without MoE layers. Finally, we demonstrate that a Rail-only interconnect only incurs 8.2%–11.2% throughput overhead for LLMs with MoEs that require all-to-all traffic.
@@ II-A Intra-platform Connectivity: High-bandwidth Domain
For instance, Nvidia’s DGX H100 server [14] consists of eight H100 GPUs interconnected with NVSwitches, providing 3.6 Tbps of non-blocking bandwidth internally. The GB200 NVL72 computer announced recently connects 36 GB200 Superchips with fifth-generation NVLink within a rack at 7.2 Tbps per GPU [15]. The AMD MI300X platform, on the other hand, employs AMD’s Infinity Fabric to connect eight MI300X accelerators in a full-mesh topology with 3.6 Tbps of bandwidth per GPU [6]. Similar platforms such as Nvidia’s DGX GH200 Super Computer have utilized multi-tiered NVSwich topologies to scale the platform’s size up to 256 GPUs while maintaining 3.6 Tbps full-bisection intra-GPU bandwidth [16]. This paper refers to a platform with Tbps internal bandwidth connectivity as a “high-bandwidth (HB) domain”, and the corresponding interconnect as HB interconnects (HBI).
@@ III-B Traffic in the NIC Domain for LLMs
Figure 4a shows the traffic pattern of training GPT-1T with hierarchical collective communication spanning both NIC and HB domains, differing from the MegatronLM cases. We compute and analyze the traffic pattern of training the GPT-1T model, with a batch size of 4096, distributed in a cluster composed of 16 Nvidia DGX GH200 supercomputers [16] (4096 GPUs). Each DGX GH200 supercomputer comprises a two-tier NVSwitch architecture, facilitating 3.6 Tbps GPU-to-GPU bandwidth across 256 H100 GPUs. Additionally, each GPU has a Connect-X7 HCA Infiniband network interface [16], which provides 400 Gbps network bandwidth in/out of each GPU. In this setup, each DGX GH200 supercomputer forms an HB domain. Figure 4 illustrates the traffic volume percentage and heatmap in this setting. The parallelization strategy has a total data parallel degree of 64, which spans 32 GPUs in each HB domain and two HB domains across the network. Figures 4b and 4c illustrate the traffic heatmap of the hierarchical AllReduce algorithm, which splits the AllReduce traffic among each DP group. Note that the network traffic stays within a rail (GPUs with a distance of 256 apart). The hierarchical algorithm efficiently utilized the bandwidth in the HB domain to carry $98\%$ of the AllReduce traffic, while the network domain carries the other $2\%$.
@@ IV-A Architecture Design
Figure 6 illustrates our Rail-only network architecture. Compared to a conventional Rail-optimized GPU cluster, shown in Figure 1, our Rail-only network keeps the HB domains but omits the full-bisection connectivity for all GPUs in the NIC domain. Instead, we only ensure that GPUs within each rail are connected with a full-bisection network.
Figure 6: Our proposal: a Rail-only network providing full-bisection connectivity to individual rails.
A straightforward illustration of our proposed architecture is to remove the spine switches (Figure 1) and re-purpose the uplinks connecting rail switches to the spine as downlinks to GPUs. Hence, a dedicated Clos network connects each rail. Compared to the Rail-optimized architecture, the Rail-only design saves the number of switches and transceivers by building multiple smaller Clos networks for individual rails, requiring fewer layers of switches in the network.
@@ IV-B Routing in Rail-only Networks
The two terms differ by ${y(x-1)D}/{C_{F}}$, which is the cost of forwarding within HB domains. When $y(x-1)\approx x(y-1)$, the slow-down factor is approximately $C_{S}/C_{F}$, which equals to $8.2\%$ and $11.2\%$ for the DGX A100 and DGX H100 generations of GPU platforms, respectively. This factor is low because of the bandwidth asymmetry between the two domains. Furthermore, this slow-down only applies to the all-to-all communication, which accounts for $27\%$ of the total traffic as shown in Figure 5. Therefore, this small overhead is negligible by Amdah’s law. We note that such properties also make our network design suitable for other classes of DNN models.
@@ V-A Iteration Time Modeling
Figure 7 compares the HFUs approximated by our analytical model with the ground truth for different GPT models and cluster scales. For GPT-1T, our computed HFU only differs from the ground truth by $1.8\%$. The discrepancy between our analytical model and the ground truth comes from our idealistic modeling of GPU computation and communication, assumptions on how computation and communication overlap, and underestimation of memory overhead. Such discrepancy goes up as the model size decreases. For the GPT-22B model, our estimation error is $10.8\%$ compared to ground truth MFU. The rest of this section utilizes our analytical model to estimate the training iteration time of Rail-only interconnects.
@@ V-B What is the Optimal Size of an HB domain?
As depicted in Figure 8, iteration times decrease as the HB domain size increases, indicating that larger HB domains reduce the network overhead during training. In both GPT models, the iteration time achieved with an HB domain size of 256 is nearly optimal. Compared to the ideal case (all GPUs are under a monolithic HB domain), GPT-146B with an HB domain of 256 is $4.1\%$ slower, while GPT-1T is $0.9\%$ slower. However, the marginal gain decreases as the HB domain size increases. For the larger GPT-1T model, the train iteration time plateaus above 32 GPUs due to Amdhahl’s law, suggesting diminishing returns from the increased cost of bigger HB domains.
@@
For a smaller GPT-146B model, shown in Figure 8b, the marginal gain of increasing HB domain size is higher than that of GPT-1T. Providing an HB domain of size eight reduces the iteration time by $43.3\%$ compared to the HB domain of size one, while increasing the HB domain size from 8 to 256 further achieves a $30.6\%$ reduction in iteration time. The more significant marginal gain for smaller LLMs incurs more communication overhead when distributed to the same cluster than larger models. This effect arises from how computation and communication costs scale as LLM grows. The communication requirement increases linearly with the model’s hidden size and sequence length. On the other hand, the model FLOPs increase quadratically with these two parameters, as indicated by previous work [21]. Therefore, we conclude that the optimal HB domain size depends on the size of the GPT model.
@@ V-D Impact of Batch Size on Network Design
While the batch size is typically an ML-centric metric for optimization, our analysis indicates that the impact of batch size on the total training time goes beyond the total number of iterations required for convergence. To further understand such impact, we analyze the iteration time of a GPT-1T model on a 32768 GPU cluster while changing the HB domain size from $K=8$ to $32768$. In this study, we vary the global batch size from 256 to 4096. Figure 10a plots the change in iteration time as the batch size varies. The iteration time exhibits a similar trajectory for all HB domain sizes; however, the relative performance (the ratio of the iteration time for an HB domain size to that of the ideal case) improves as the batch size increases. Figure 10b represents this trend. When $K=256$, the relative performance increases from $95\%$ to $99\%$ as the batch size increases from 256 to 4096 sequences. The iteration time advantage is prominent when the HB domain size is small. For $K=8$, increasing the batch size from 256 to 4096 improves the relative performance from $65\%$ to $85\%$, suggesting a larger batch size is preferable for a cluster with a smaller HB domain. Prior studies have shown that LLM training benefits from a larger batch size [29, 1], especially for bigger models, making it a perfect fit for our Rail-only design.